OpenStudy (anonymous):

what is the antiderivative of (x^2)/(9 + x^(6))

OpenStudy (amistre64):

\[\int \frac{x^2}{9+x^6}dx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You need to apply integration by parts twice to the function \[x^2(9+x^6)^{-1}\]

OpenStudy (amistre64):

that would work out great :)

OpenStudy (anonymous):

Or maybe not (how to find integral of 1/9+x^6?

OpenStudy (amistre64):

..... if we could do that

OpenStudy (amistre64):

\[(x^2)^3\] comes to mind, but may not be helpful

OpenStudy (anonymous):

Substitute y = x^3... it should work I think!

OpenStudy (amistre64):

\[(3x^3)^2\]would be nice maybe

OpenStudy (anonymous):

use the substitution u=x^3? gives you arctan(x^3/3)/9??

OpenStudy (anonymous):

Possible intermediate steps: integral x^2/(9+x^6) dx For the integrand x^2/(x^6+9), substitute u = x^3 and du = 3 x^2 dx: = 1/3 integral 1/(u^2+9) du The integral of 1/(u^2+9) is 1/3 tan^(-1)(u/3): = 1/9 tan^(-1)(u/3)+constant Substitute back for u = x^3: = 1/9 tan^(-1)(x^3/3)+constant

OpenStudy (anonymous):

Let y = x^3 dy/dx = 3x^2 dy/3 = x^2 dx \[\int{\frac{x^2}{9+x^6}}dx=1/3\int{\frac{1}{9+y^2}}dy\] which can be evaluated

OpenStudy (amistre64):

straight off the wolfram lol

OpenStudy (anonymous):

yeh hate that dont you amistre

OpenStudy (amistre64):

oh i like it when i get stuck, but to claim it as your own is a bit misleading ;)

OpenStudy (anonymous):

yeah, substituting y=x^3 works a treat :)

OpenStudy (anonymous):

yeh thats what i mean haha its a godsend when your stuck

OpenStudy (amistre64):

Wolfram gives us: Possible intermediate steps: integral x^2/(9+x^6) dx For the integrand x^2/(x^6+9), substitute u = x^3 and du = 3 x^2 dx: = 1/3 integral 1/(u^2+9) du The integral of 1/(u^2+9) is 1/3 tan^(-1)(u/3): = 1/9 tan^(-1)(u/3)+constant Substitute back for u = x^3: = 1/9 tan^(-1)(x^3/3)+constant

OpenStudy (anonymous):

I got the answer already guys :)

OpenStudy (anonymous):

so then from there xactxx where do i go?

OpenStudy (amistre64):

... on to the next problem :)

OpenStudy (anonymous):

hahahaah good answer :)

OpenStudy (anonymous):

arctangent yes?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

yessss